# Astute RAG: Overcoming Imperfect Retrieval Augmentation and Knowledge Conflicts for Large Language Models

### v1

@@ Abstract
Retrieval-Augmented Generation (RAG), while effective in integrating external knowledge to address the limitations of large language models (LLMs), can be undermined by imperfect retrieval, which may introduce irrelevant, misleading, or even malicious information. Despite its importance, previous studies have rarely explored the behavior of RAG through joint analysis on how errors from imperfect retrieval attribute and propagate, and how potential conflicts arise between the LLMs’ internal knowledge and external sources. We find that imperfect retrieval augmentation might be inevitable and quite harmful, through controlled analysis under realistic conditions. We identify the knowledge conflicts between LLM-internal and external knowledge from retrieval as a bottleneck to overcome in the post-retrieval stage of RAG. To render LLMs resilient to imperfect retrieval, we propose Astute RAG, a novel RAG approach that adaptively elicits essential information from LLMs’ internal knowledge, iteratively consolidates internal and external knowledge with source-awareness, and finalizes the answer according to information reliability. Our experiments using Gemini and Claude demonstrate that Astute RAG significantly outperforms previous robustness-enhanced RAG methods. Notably, Astute RAG is the only approach that matches or exceeds the performance of LLMs without RAG under worst-case scenarios. Further analysis reveals that Astute RAG effectively resolves knowledge conflicts, improving the reliability and trustworthiness of RAG systems.
@@ 1 Introduction
Retrieval augmented generation (RAG) has become the standard approach for large language models (LLMs) to tackle knowledge-intensive tasks (Guu et al., 2020; Lewis et al., 2020). Prior works mainly leverage RAG to address the inherent knowledge limitations of LLMs, effectively integrating missing information and grounding to reliable sources. However, recent research has highlighted a significant drawback that RAG might rely on imperfect retrieval results, including irrelevant, misleading, or even malicious information, which eventually leads to inaccurate LLM responses (Chen et al., 2024a; Xiang et al., 2024; Zou et al., 2024). For example, when asked about the practice of eating rocks, LLMs might cite misleading information, such as a satirical news source claiming that one should consume at least one rock per day. [^1] https://www.bbc.com/news/articles/cd11gzejgz4o. The occurrence of imperfect retrieval augmentation is inevitable, driven by factors such as corpus quality limitations (Shao et al., 2024), the reliability of retrievers (Dai et al., 2024), and the complexity of the queries (Su et al., 2024). This poses a significant challenge to the trustworthiness of RAG.
While there have been independent analyses of information retrieval and RAG in the context of LLMs (Su et al., 2024; Mallen et al., 2023), previous studies have rarely connected the behaviors of retrieval and subsequent generation, particularly regarding the propagation of information retrieval errors, which may lead to knowledge conflicts (Longpre et al., 2021; Wang et al., 2023a; Xu et al., 2024b) between LLMs and context. To this end, we conduct comprehensive analyses on the occurrence of imperfect retrieval augmentation and its impact on LLM behavior under realistic conditions (Section 2). We conduct controlled experiments on a diverse range of general, domain-specific, and long-tail questions from NQ (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), BioASQ (Tsatsaronis et al., 2015), and PopQA (Mallen et al., 2023). We observe that imperfect retrieval augmentation is widespread even with adept real-world search engine (such as Google Search with Web as corpus) – roughly 70% retrieved passages do not directly contain true answers, leading to the impeded performance of LLM with RAG augmentation. [^2] Note that some passages may contain information indirectly relevant to the answer, but may unintentionally mislead or distract LLMs.
These findings underscore the potential severity of the imperfect retrieval issue in real-world RAG and highlight the widespread existence of knowledge conflicts as the bottleneck to overcome it (Figure 1). Recent studies demonstrate that LLM-internal and external knowledge offer distinct advantages, but LLMs often struggle to consolidate conflicting information reliably, failing to respond based on collective knowledge (Mallen et al., 2023; Tan et al., 2024; Xie et al., 2024; Jin et al., 2024). This raises the following research question: Is there an effective method to combine internal (from LLMs’ pretrained weights) and external (from specific corpora or knowledge bases) knowledge for more reliable RAG? Previous work has widely explored using external knowledge to enhance LLMs through RAG. We seek to further leverage LLMs’ internal knowledge to recover from RAG failures
@@
Figure 1: Knowledge conflicts between the LLMs’ internal knowledge and retrieved knowledge from external sources. We report the overall results with Claude under the setting in Section 4.1.
@@
Motivated by these important real-world challenges, we propose Astute RAG (Section 3), a novel RAG approach designed to be resilient to imperfect retrieval augmentation, while preserving RAG grounding effect when RAG is reliable. To this end, Astute RAG needs effectively differentiate the reliability of the LLM’s intrinsic knowledge and the external information retrieved in RAG, utilizing each only when trustworthy and ensuring proper integration. Specifically, Astute RAG initially elicits information from LLMs’ internal knowledge to explicitly complement the passages retrieved from external sources. Then, Astute RAG conducts source-aware knowledge consolidation of information from various internal and external sources. The desiderata is combining consistent information, identifying conflicting information, and filtering out irrelevant information. Finally, Astute RAG proposes answers based on each group of consistent passages and compares the answers from different passage groups to determine the final answer. Our experiments involving Gemini and Claude [^3] https://www.anthropic.com/claude on various datasets (Section 4) demonstrate the superior performance of Astute RAG compared to previous RAG approaches designed to be robust against retrieval corruptions. Moreover, Astute RAG consistently outperforms baselines across different retrieval quality levels. Notably, Astute RAG is the only RAG method that achieves performance comparable to or even surpassing conventional use of LLMs under the worst-case scenario where all retrieved passages are unhelpful. Further analysis reveals the effectiveness of Astute RAG in resolving knowledge conflicts between internal and external knowledge.
@@
To conclude, our core contributions are threefold. First, we analyze RAG under realistic conditions, identifying imperfect retrieval augmentation as a significant contributor to RAG failures and pinpointing knowledge conflicts as the primary bottleneck in overcoming it. Second, we propose Astute RAG, which explicitly addresses conflicts between LLM-internal and external knowledge, thereby recovering from RAG failures. Third, experiments with various LLMs and datasets demonstrate the effectiveness of Astute RAG, even in the most challenging scenarios.
2 Imperfect Retrieval: The Pitfall of RAG
@@
To better showcase the common real-world challenges and to make better motivate for improved methodological designs, we evaluate retrieval quality, end-to-end RAG performance, and knowledge conflicts on a controlled set of data. The selected data encompass a diverse range of general, domain-specific, and long-tail questions from NQ (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), BioASQ (Tsatsaronis et al., 2015), and PopQA (Mallen et al., 2023). Our analysis is based on realistic retrieval results with Google Search [^4] https://developers.google.com/custom-search/v1/overview as the retriever and the Web as the corpus. This setting allows us to analyze the severity of imperfect retrieval in real-world RAG. Overall, we sample 1K short-form QA instances from these datasets, and pair each instance with 10 retrieved passages.
Imperfect retrieval is common. We examine the occurrence of correct answers in retrieved passages as an approximation of retrieval quality. Since we mainly focus on short-form QA which provides most variants of the correct answer for each question, the approximation through string matching can give us a rouge intuition of how precise the retrieval result is. Specifically, we define the retrieval precision as the ratio of passages containing the correct answer for each instance:
Retrieval Precision={number of retrieved passages containing correct answer}{number of total retrieved $\text{Retrieval Precision}=\frac{\text{\{number of retrieved passages containing correct answer\}}}{\text{\{number of total retrieved passages\}}}$
As shown in Figure 2, although instances from different datasets exhibit different data distributions, imperfect retrieval is prevalent. Specifically, $\sim$20of the overall data have no mentions of the correct answer within any retrieved passage, including 34% on NQ, 18% on TriviaQA, 24% on BioASQ, and 50% on PopQA. This finding also aligns with previous observation on information retrieval (Thakur et al., 2024), that highlights that the number of positive passages can be very limited.
Imperfect retrieval leads to RAG failures. We further analyze the relation between retrieval quality and RAG performance. We compare the performance of Claude 3.5 Sonnet, with and without RAG and report the results by retrieval precision in Figure 4. In general, RAG is helpful when the retrieval precision is not lower than 20%. When the retrieval precision is close to 0, the model with RAG performs much worse than without RAG, indicating that imperfect retrieval augmentation can be the cause of RAG failures. This finding aligns with the previous observation from Yu et al. (2024) that adding more retrieved passages does not necessarily lead to better performance, as the additional passages might reduce the retrieval precision.
Knowledge conflicts widely exist in RAG failures. We provide an in-depth analyses of knowledge conflicts between LLMs’ internal knowledge and retrieved passages from external sources. With Claude 3.5 Sonnet as the LLM, Figure 1 shows that 19.2% of the overall data exhibit knowledge conflicts, where either the answer with or without RAG is correct. Among the conflicting cases, the internal knowledge is correct on 47.4% of them, while the external knowledge is correct on the remaining 52.6%. These results emphasize the importance of effectively combining the internal and external knowledge to overcome the inherent limitation of relying solely on either source. However, previous work (Tan et al., 2024; Xie et al., 2024; Jin et al., 2024) show that LLMs might respond based on misleading information rather than comprehensive understanding of the conflicting knowledge in this context.
3 Astute RAG: Overcoming the Pitfall
@@
Figure 3: Overview of the proposed Astute RAG framework. Astute RAG is designed to better combine the information from the external sources (e.g. web, domain-specific corpora, knowledge bases) and internal knowledge of the LLMs by employing a consolidation mechanism to address the information conflicts, which eventually leads to better quality generated outputs.
We begin with formulating the problem of imperfect retrieval in RAG (Section 3.1). We then provide an overview of Astute RAG, designed to overcome this problem (Section 3.2). Subsequently, we delve into the three major steps of Astute RAG, including adaptive generation of internal knowledge (Section 3.3), source-aware knowledge consolidation (Section 3.4), and answer finalization (Section 3.5).
3.1 Problem Formulation
Our objective is to mitigate the effects of imperfect retrieval augmentation, resolve knowledge conflicts between the LLM’s internal knowledge and external sources (such as custom/public corpora and knowledge bases), and ultimately produce more accurate and reliable responses from LLMs.
Given a set of retrieved passages from external sources $E=[e_{1},\ldots,e_{n}]$, a pre-trained LLM $\mathcal{M}$ (accessible through prediction-only APIs, encompassing commercial black-box ones), and a query $q$, the task is to generate the corresponding correct answer $a^{*}$. Notably, this setting is orthogonal to prior work on improving the retriever, training LLMs, or conducting adaptive retrieval, which are mainly preliminary steps.
3.2 Overview of the Framework
Astute RAG is designed to better leverage collective knowledge from both internal knowledge of LLMs and external corpus, for more reliable responses. As shown in Figure 3 and Algorithm 1, Astute RAG starts from acquiring the most accurate, relevant, and thorough passage set from the LLMs’ internal knowledge. Then, internal and external knowledge are consolidated in an iterative way, by comparing the generated and retrieved passages. Finally, the reliability of conflicting information is compared and the final output is generated according to the most reliable knowledge.
@@ 3.2 Overview of the Framework
2:Adaptively generate passages: $I\leftarrow\mathcal{M}(p_{gen},q,\hat{m})$triangleright Section 3.3
@@
9:     Finally consolidate and answer: $a\leftarrow\mathcal{M}(p_{ans},q,\langle D_{0},S_{0}\rangle,\langle D_{t-1},S_{t-1}\rangle)$triangleright Section 3.5
@@ 3.3 Adaptive Generation of Internal Knowledge
Inspired by Constitutional AI (Bai et al., 2022), we provide constitutional principles indicating the desired properties of internal passages in the prompt $p_{gen}$ (see Appendix A for details) to guide their generation, emphasizing that the generated passages should be accurate, relevant, and hallucination-free. Moreover, we allow the LLM to perform adaptive generation of passages in its internal knowledge. The LLM can decide how many passages to generate by itself. Rather generating a fix number of passages, we request the LLM to generate at most $\hat{m}$ passages, each covering distinct information, and to directly indicate if no more reliable information is available. This adaptive approach allows the LLM to generate fewer passages (or even no passages at all) when the useful information within internal knowledge is limited and more passages when there are multiple feasible answers in the internal knowledge. In this step, the LLM generates $m\leq\hat{m}$ passages based on its internal knowledge:

### v2

@@ Abstract
Retrieval-augmented generation (RAG), while effective in integrating external knowledge to enhance large language models (LLMs), can be undermined by imperfect retrieval, which may introduce irrelevant, misleading, or even malicious information. Despite its importance, previous studies have rarely explored the behavior of RAG with errors from imperfect retrieval, and how potential conflicts arise between the LLMs’ internal knowledge and external sources. We show that imperfect retrieval augmentation might be inevitable and quite harmful, through controlled analysis under realistic conditions. Knowledge conflicts between LLM-internal and external knowledge from retrieval is a bottleneck to overcome in the post-retrieval stage of RAG. To render LLMs resilient to imperfect retrieval, we propose Astute RAG, a novel RAG approach that adaptively elicits essential information from LLMs’ internal knowledge, iteratively consolidates internal and external knowledge with source-awareness, and finalizes the answer according to information reliability. Our experiments with Gemini and Claude demonstrate that Astute RAG significantly outperforms previous robustness-enhanced RAG methods. Notably, Astute RAG is the only approach that matches or exceeds the performance of LLMs without RAG under worst-case scenarios. Astute RAG effectively resolves knowledge conflicts, improving the reliability and trustworthiness of RAG systems.
Astute RAG: Overcoming Imperfect Retrieval Augmentation and Knowledge Conflicts for Large Language Models
Fei Wang12∗  Xingchen Wan1  Ruoxi Sun1  Jiefeng Chen1  Sercan Ö. Arık1 1Google Cloud AI Research  2University of Southern California fwang598@usc.edu{xingchenw,ruoxis,jiefengc,soarik}@google.com
11footnotetext: Work done during internship at Google.
@@ 1 Introduction
Retrieval-augmented generation (RAG) is commonly used for large language models (LLMs) to tackle knowledge-intensive tasks (Guu et al., 2020; Lewis et al., 2020). Prior works mainly leverage RAG to address the inherent knowledge limitations of LLMs, effectively integrating missing information and grounding to reliable sources. However, recent research has highlighted a significant drawback that RAG might rely on imperfect retrieval, including irrelevant, misleading, or even malicious information (Fig. 1), which eventually leads to inaccurate LLM responses (Chen et al., 2024a; Zou et al., 2024). Moreover, recent studies have shown that retrieval augmentation can confuse LLMs when retrieved passages are conflicting with LLMs’ parametric knowledge Tan et al. (2024); Xie et al. (2024); Jin et al. (2024). These pose significant challenges to the trustworthiness of RAG.
@@
Figure 1: Knowledge conflicts between the LLMs’ internal knowledge and retrieved knowledge from external sources. We report the overall results with Claude under the setting in Sec. 5.1.
To address imperfect retrieval, earlier work seeks to improve the retrieval approaches, such as dynamic and iterative retrieval Jiang et al. (2023); Asai et al. (2023); Yan et al. (2024). However, the occurrence of imperfect retrieval is still inevitable, due to corpus quality limitations (Shao et al., 2024), the reliability of retrievers (Dai et al., 2024), and the complexity of queries (Su et al., 2024). Consequently, recent work shifts the focus to the generation stage, seeking to reduce the negative impact of noisy retrieved passages Xiang et al. (2024); Wei et al. (2024). Another line of research at generation stage, motivated by knowledge conflicts, has explored complementing retrieved passages with LLM-generated passages Yu et al. (2023a); Zhang et al. (2023) or deactivating RAG when the retrieved passages are of insufficient quality Xu et al. (2024); Mallen et al. (2023); Jeong et al. (2024).
Despite the previous work on the impact of imperfect retrieval and knowledge conflicts at RAG generation stage, quantitative analyses lack on two crucial real-world aspects: (i) the relation between retrieval quality and occurrence of knowledge conflicts, and (ii) the extent to which retrieved passages and LLMs’ parametric knowledge can correct each other. Method-wise, existing approaches for mitigating RAG failures caused by imperfect retrieval and knowledge conflicts have not yet yielded a training-free method capable of explicitly analyzing conflicting knowledge across various internal and external sources, and achieving worst-case robustness for black-box LLMs.
In this paper, we first conduct comprehensive analyses to investigate the relation between imperfect retrieval and knowledge conflicts, and examine the frequency of external and LLMs’ internal knowledge mutually correcting each other (Sec. 3). On a diverse range of general, domain-specific, and long-tail questions from NQ (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), BioASQ (Tsatsaronis et al., 2015), and PopQA (Mallen et al., 2023), we observe that imperfect retrieval is widespread even with an adept real-world search engine, leading to the impeded performance of RAG.111such as Google Search with Web as corpus Retrieval precision is tightly correlated with the knowledge conflict rate. Mutual correction between the LLM’s knowledge and external knowledge is crucial for recovering from RAG failures. Our findings underscore the potential severity of imperfect retrieval in real-world RAG and highlight the widespread existence of knowledge conflicts as the bottleneck.
@@
We propose Astute RAG, a novel RAG approach designed for resilience to imperfect retrieval augmentation, while preserving RAG grounding effect when retrieval is reliable (Sec. 4). Astute RAG effectively differentiates between consistent and conflicting information from the LLM’s internal knowledge and the externally retrieved passages, assesses their reliability, and ensures proper integration of trustworthy information. Astute RAG first adaptively elicits LLMs’ knowledge and then conducts source-aware knowledge consolidation. The desiderata is combining consistent information, identifying conflicting information, and filtering out irrelevant information. Finally, Astute RAG proposes answers based on consistent information and compares them to determine the final answer. Our experiments with various LLMs (Claude, Gemini and Mistral), demonstrate superior performance of Astute RAG compared to previous RAG approaches designed for robustness (Sec. 5). Astute RAG consistently outperforms baselines across different retrieval quality levels. Notably, Astute RAG is the only RAG method that achieves performance comparable to or even surpassing retrieval-free mode of LLMs under the worst-case scenario where all retrieved passages are unhelpful. Further analysis reveals the effectiveness of Astute RAG in resolving knowledge conflicts.
@@
In summary, our core contributions are threefold. First, we provide quantitative analyses and novel insights for the connection among imperfect retrieval, knowledge conflicts, and RAG failures under real-world conditions. Second, we propose Astute RAG, which explicitly analyzes LLM-internal and external knowledge in-context, assesses their reliablity, and recovers from RAG failures with black-box access. Third, with experiments with various LLMs and datasets, we demonstrate the effectiveness of Astute RAG in improving robustness and trustworthiness, even in the most challenging scenarios.
2 Related Work
@@
Figure 3: Conflicting rate between answers from LLMs with and without RAG on different retrieval precision.
RAG aims to address the inherent knowledge limitation of LLMs with passages retrieved from external sources of information such as private corpora or public knowledge bases (Guu et al., 2020; Lewis et al., 2020; Borgeaud et al., 2022). Given the widespread real-world adoption of RAG, including risk-sensitive domains, the negative impact of noisy information within retrieved passages has garnered increasing attention (Cuconasu et al., 2024). Recent work explored enhancing the robustness of RAG systems against noise from various perspectives, including training LLMs with noisy context (Yu et al., 2023b; Yoran et al., 2024; Pan et al., 2024; Fang et al., 2024), training small models to filter out irrelevant passages (Wang et al., 2023b; Xu et al., 2023), passage reranking (Yu et al., 2024; Glass et al., 2022), dynamic and iterative retrieval (Jiang et al., 2023; Asai et al., 2023; Yan et al., 2024), query rewriting (Ma et al., 2023), and speculative drafting (Wang et al., 2024). These focus on distinct modules or stages of RAG systems and are orthogonal to our work.
Our work focuses on enhancing RAG robustness at the post-retrieval stage, after retrieved passages have been provided. On this, RobustRAG (Xiang et al., 2024) aggregates answers from each independent passage to provide certifiable robustness. InstructRAG (Wei et al., 2024) instructs the LLM to provide a rationale connecting the answer with information in passages. MADRA (Wang et al., 2023a) applies multi-agent debate to select helpful evidence. However, these do not explicitly incorporate internal knowledge to recover from RAG failures and therefore might severely suffer when the majority of retrieved passages have issues. For emphasizing internal knowledge of LLMs in RAG, recent work explored using LLM-generated passage as context (Yu et al., 2023a), training models to match generated and retrieved passages (Zhang et al., 2023), adaptively switching between LLMs with and without RAG (Xu et al., 2024; Mallen et al., 2023; Jeong et al., 2024), and combining answers through contrastive decoding (Zhao et al., 2024; Jin et al., 2024). Different from prior work, we provide a systematic framework on connecting imperfect retrieval, knowledge conflicts, and RAG failures. Specifically focusing on the imperfect context setting, our method is training-free and applicable to black-box LLMs, explicitly analyzes internal and external knowledge in-context, and offers broader usability and adaptability.
3 The Pitfall of RAG
@@
To better showcase common real-world challenges and motivate improved methodological designs, we evaluate retrieval quality, the occurrence of knowledge conflicts, their relationship, and the mutual correction between external and internal knowledge using a controlled dataset derived from NQ, TriviaQA, BioASQ, and PopQA, datasets widely used for RAG in prior work Xiang et al. (2024); Wei et al. (2024); Asai et al. (2023). Different from prior work, our analysis is based on real-world retrieval results with Google Search [^2] https://developers.google.com/custom-search/v1/overview as the retriever and the Web as the corpus. Overall, we sample 1K instances, each with 10 retrieved passages.
@@
Figure 4: Overview of the Astute RAG framework. Astute RAG is designed to better combine the information from the external sources (e.g. web, domain-specific corpora) and internal knowledge of the LLMs by employing a consolidation mechanism to address the conflicts, which eventually leads to superior generation quality.
Imperfect retrieval and knowledge conflicts are common and harmful. Our initial observations are consist with prior work. As shown in Fig. 2, the retrieval precision [^3] Ratio of passages directly contain true answers. is generally low - roughly 70% retrieved passages do not directly contain true answers, consistent with prior work demonstrating the often imperfect nature of retrieval results Thakur et al. (2024); Su et al. (2024). With Claude 3.5 Sonnet as the LLM, Fig. 1 shows that 19.2% of the overall data exhibit knowledge conflicts, consistent with prior work demonstrating the prevalence of such conflicts across various scenarios Pham et al. (2024); Xie et al. (2024); Longpre et al. (2021). Moreover, we observe strong correlations between retrieval precision and RAG performance (Fig. 7) and between the occurrence of knowledge conflicts and RAG performance (Fig. 8), findings consistent with prior work on these respective topics Chen et al. (2024a); Xie et al. (2024).
Lower retrieval precision increases knowledge conflicts in general. As shown in Fig. 3, most advanced LLMs exhibit the highest conflict rates when retrieval precision is as low as 10%. Subsequently, the conflict rate generally decreases as precision increases, although some fluctuations may occur. This trend is generally applicable to the studied LLMs with different training processes. Notably, when retrieval precision is 0%, conflict rates tend to be significantly lower. This suggests that limited external knowledge for the query results in more irrelevant passages rather than incorrect ones.
Internal and external knowledge can correct each other to a comparable extent. Among the conflicting cases, the internal knowledge is correct on 47.4% of them, while the external knowledge is correct on the remaining 52.6%. These results emphasize the importance of effectively combining the internal and external knowledge to overcome the inherent limitation of relying solely on either source. However, previous work (Tan et al., 2024; Xie et al., 2024; Jin et al., 2024) shows that LLMs often select knowledge based on unreliable shortcuts, so simply presenting LLM-generated passages in the context may not help.
4 Astute RAG
@@
2:Adaptively generate passages: $I\leftarrow\mathcal{M}(p_{gen},q,\hat{m})$triangleright Sec. 4.2
@@
9:     Finally consolidate and answer: $a\leftarrow\mathcal{M}(p_{ans},q,\langle D_{0},S_{0}\rangle,\langle D_{t-1},S_{t-1}\rangle)$triangleright Sec. 4.4
@@
We first provide an overview of Astute RAG (Sec. 4.1). Subsequently, we delve into the three major steps of Astute RAG, including adaptive generation of internal knowledge (Sec. 4.2), source-aware knowledge consolidation (Sec. 4.3), and answer finalization (Sec. 4.4).
4.1 Overview
Our objective is to mitigate the effects of imperfect retrieval augmentation, resolve knowledge conflicts between the LLM’s internal knowledge and external sources (such as custom/public corpora and knowledge bases), and ultimately produce more accurate and reliable responses from LLMs. Given a set of retrieved passages from external sources $E=[e_{1},\ldots,e_{n}]$, a pre-trained LLM $\mathcal{M}$ (accessible through prediction-only APIs, encompassing commercial black-box ones), and a query $q$, the task is to generate the corresponding correct answer $a^{*}$. Notably, this setting is orthogonal to prior work on improving the retriever, training LLMs, or conducting adaptive retrieval, which are mainly preliminary steps.
Astute RAG is designed to better leverage collective knowledge from both internal knowledge of LLMs and external corpus, for more reliable responses. As shown in Fig. 4 and Alg. 1, Astute RAG starts from acquiring the most accurate, relevant, and thorough passage set from the LLMs’ internal knowledge. Then, internal and external knowledge are consolidated in an iterative way, by comparing the generated and retrieved passages. Finally, the reliability of conflicting information is compared and the final output is generated according to the most reliable knowledge.
@@ 4.2 Adaptive Generation of Internal Knowledge
Inspired by Bai et al. (2022), we provide constitutional principles indicating the desired properties of internal passages in the prompt $p_{gen}$ (see Appx. A for details) to guide their generation, emphasizing that the generated passages should be accurate, relevant, and hallucination-free. Moreover, we allow the LLM to perform adaptive generation of passages in its internal knowledge. The LLM can decide how many passages to generate by itself. Rather generating a fix number of passages, we request the LLM to generate at most $\hat{m}$ passages, each covering distinct information, and to directly indicate if no more reliable information is available. This adaptive approach allows the LLM to generate fewer passages (or even no passages at all) when the useful information within internal knowledge is limited and more passages when there are multiple feasible answers in the internal knowledge. In this step, the LLM generates $m\leq\hat{m}$ passages based on its internal knowledge: